OSMOS Mechanism Numbering
Inside the IEB
Here is a schematic of the inside of the instrument electronics box with
the MicroLYNX numbered.
Outside the IEB
Here is how the numbering corresponds to the mechanism cable positions on the
IEB panel. These numbers correspond to the individual MicroLYNX. Mechanism
cables can be swapped as necessary to the open position if a MicroLYNX fails.
The code for that mechanism must be uploaded into the spare MicroLYNX.
Here is how the positions on the OSMOS panel correspond to the mechanisms
within the instrument. This does not change.
The as-shipped configuration was as follows:
|1||8001||Slit Wheel||slit||2||8002||Collimator Focus||colfoc
||5||8005||Filter Wheel 1||filter1
||7||8007||Filter Wheel 2||filter2
The current configuration (as of 2017 Aug 24) is:
|1||8001||Slit Wheel||slit||2015 Feb 5||rwp/osu||2||8002||Collimator Focus||colfoc||2014 Feb 7||rwp/osu
||4||8004||Camera Focus||camfoc||2012 Oct||rwp/osu
||5||8005||Disperser||disperser||2015 Oct 13||rwp/osu
||6||8006||Filter Wheel 1||filter1||2015 Feb 5||rwp/osu
||7||8007||Filter Wheel 2||filter2||2014 Feb 7||rwp/osu
Trouble Record by MicroLYNX position
The large number of failures and flaky behavior in October 2012
followed a thunderstorm with frequent cloud-to-ground lightning near
the observatory, so the probably cause was lightning surge damage.
- 2012 Oct: Replaced with new MicroLYNX Oct 2012, started showing intermittent failures in May 2012, but no outright failures or the red fault LED light.
- 2014 Feb: replaced unit, original showing flaky behavior.
- 2015 Dec 28: Failed, showing red LED indicator, swapped disperser to MLC5
- 2012 Oct: showing flaky behavior, replaced with new unit
- 2012 Oct: Failed (red LED), replaced with new unit.
- 2015 Feb 5: Replaced failed unit
- 2015 Feb 5: Replaced Failed unit.
- 2014 Feb: Replaced after orginal failed Jan 2014
Additional surge protection circuitry was introduced in 2015 following
troubleshooting of failed/flaky units to OSU for testing.
Updated: 2017 Aug 24 [rwp/osu]